# An explicit form of Ingham’s zero density estimate

### v1

@@ 4. Proof of Theorem 1
$F_{X}\left(\sigma,H,T,w\right)\leq\left(\frac{\exp{\left(\kappa\left(1+\left(\frac{1}{4T}\left(1+\frac{2\delta}{\log{(hT)}}\right)\right)^{2}\right)\right)}}{1-H^{-1}}\right)^{\frac{1}{w}}\times\\ \times\left(\left(K_{1}^{\frac{\frac{3}{4}w_{1}}{w}}K_{3}^{\frac{\frac{1}{2}w_{2}}{w}}\right)T^{\frac{\frac{3}{4}w_{1}}{w}}\left(\log{T}\right)^{\frac{w_{1}}{w}}\left(\log{(hT)}\right)^{\frac{\frac{1}{2}w_{1}+w_{2}}{w}}+\left(K_{2}^{\frac{\frac{3}{4}w_{1}}{w}}K_{3}^{\frac{\frac{1}{2}w_{2}}{w}}\right)T^{\frac{\frac{3}{4}w_{1}}{w}}\left(\log{(hT)}\right)^{\frac{w_{2}}{w}}\right)$ (42)
@@
$K_{3}^{\frac{\frac{1}{2}w_{2}}{w}}\leq\max\left\{1,K_{3}^{-\frac{3\delta}{\log{(hT_{0})}}}\right\}K_{3}^{\frac{2\sigma-1}{2-\sigma}}.$ (45)
Turning to the log-terms, because $\log{T}\geq 1$ and
@@
$\left(\log{T}\right)^{\frac{w_{1}}{w}}\leq\exp{\left(\frac{4\delta\log{\log{T}}}{\log{(hT)}}\right)}\left(\log{T}\right)^{\frac{4(1-\sigma)}{2-\sigma}}\leq\exp{\left(4\delta\max\left\{\frac{\log{\log{T_{0}}}}{\log{(hT_{0})}},\frac{\log{\log{T_{0}}}}{\log{T_{0}}}\right\}\right)}\left(\log{T}\right)^{\frac{4(1-\sigma)}{2-\sigma}},$
where we also used $T_{0}\geq e^{e}$. Moreover,
$\frac{\frac{1}{2}w_{1}+w_{2}}{w}=\frac{2\sigma+\frac{2\delta}{\log{(hT)}}}{2-\sigma+\frac{3\delta}{\log{(hT)}}}\leq\frac{2\sigma}{2-\sigma},\quad\frac{w_{2}}{w}=\frac{2(2\sigma-1)}{2-\sigma+\frac{3\delta}{\log(hT)}}\leq\frac{2(2\sigma-1)}{2-\sigma},$
@@
$\left(\log{(hT)}\right)^{\frac{\frac{1}{2}w_{1}+w_{2}}{w}}\leq\left(\log{(hT)}\right)^{\frac{2\sigma}{2-\sigma}},\quad\left(\log{(hT)}\right)^{\frac{w_{2}}{w}}\leq\left(\log{(hT)}\right)^{\frac{2(2\sigma-1)}{2-\sigma}}$ (46)
since $\log{(hT)}\geq 1$. Finally,
$\frac{1}{w}=\frac{2+\frac{4\delta}{\log{(hT)}}}{2-\sigma+\frac{3\delta}{\log{(hT)}}}\leq\frac{2}{2-\sigma}$
is used to bound the first term on the right-hand side of (42), and this contributes $L_{0}$ from (35) to the final estimate. Inequalities (43)–(46) contribute to $L_{1}$ from (39), while inequality (43) for $K_{2}$ in place of $K_{1}$, and inequalities (44), (45) and (46), contribute to $L_{2}$ from (40). The proof of Lemma 9 is thus complete. ∎
@@
×$\times\left(L_{1}\max\left\{1,K_{1}^{\frac{3d}{\log{T_{0}}}}\right\}\exp{\left(d\left(3+\frac{4\log{\log{T_{0}}}}{\log{T_{0}}}\right)\right)}\left(K_{1}T\right)^{\frac{3(1-\sigma)}{2-\sigma}}\left(\log{T}\right)^{\frac{4(1-\sigma)}{2-\sigma}}\left(\log{(hT)}\right)^{\frac{2\sigma}{2-\sigma}}\right.$
$\left.+L_{2}e^{3d}\max\left\{1,K_{2}^{\frac{3d}{\log{T_{0}}}}\right\}\left(K_{2}T\right)^{\frac{3(1-\sigma)}{2-\sigma}}\left(\log{(hT)}\right)^{\frac{2(2\sigma-1)}{2-\sigma}}\right),$
where $L_{0}$, $L_{1}$, $L_{2}$ and $K_{1}$, $K_{2}$, $K_{3}$ are as in Lemma 9. It is not hard to see that the statement of Theorem 1 now follows after combining all estimates. Also, the non-vanishing condition can be removed by evoking the standard perturbation argument since the right-hand side of the main estimate of Theorem 1 is a continuous function in $\sigma$, $T$ and $H$. ∎
@@ 5.1. The method
This is a meromorphic function on $\mathbb{C}$ with only simple poles at $s=2k-1$ for $k\in\mathbb{N}$. Also note that the functional equation implies $\chi(1/2+\mathrm{i}t)\chi(1/2-\mathrm{i}t)=\left|\chi(1/2\pm\mathrm{i}t)\right|^{2}=1$. The starting point in Ramachandra’s proof (which itself resembles the reflection principle [Ivi03, Section 4.4] or a half-way derivation of a weighted approximate functional equation [Ivi13, Section 4.5]) is to use the Mellin integral for $e^{-x}$ to get
@@
The curly brackets in (52) and (53) indicate the plan of the proof of Theorem 2. Precisely, the seemingly unnatural appearance of the second and the third part allows us to obtain an asymptotically correct estimate (12) since we can show that all terms, except the first one, are $\ll T\log^{3}{T}$. However, if interested in obtaining just an upper bound like (13), then  (52) and (53) can be simplified to
@@ 5.2. Some results on various sums that contain $d(n)$ and $d^{2}(n)$
by partial summation and Lemma 13. Repeated use of partial integration gives
@@
$-e^{-2}x^{2}\sum_{k=1}^{4}D_{k}(\log{x})^{4-k}+\sum_{k=1}^{4}D_{k}\int_{x}^{\infty}\left(\frac{2}{x}u^{2}-u\right)e^{-\frac{2u}{x}}(\log{u})^{4-k}\mathrm{d}u\leq\frac{3D_{1}}{4e^{2}}x^{2}(\log{x})^{3}\\ +\left(\frac{57D_{1}}{16e^{2}}+\frac{3D_{2}}{4e^{2}}\right)x^{2}(\log{x})^{2}+\left(\frac{3D_{1}}{e^{2}}+\frac{19D_{2}}{8e^{2}}+\frac{3D_{3}}{4e^{2}}\right)x^{2}\log{x}+\left(\frac{3D_{1}}{4e^{2}}+\frac{D_{2}}{e^{2}}+\frac{19D_{3}}{16e^{2}}+\frac{3D_{4}}{4e^{2}}\right)x^{2}.$ (71)
@@ 5.3. Proof of Theorem 2
with the last inequalities true by [DZA22, Proposition 2.2]. For the second part, one can use (73) to obtain $|\chi(z)|\leq 0.4|y|^{1/2-x}$ for $x\in[-1/2,0]$ and $|y|\geq 10^{2}$, while the maximum principle and numerical verification show that $|\chi(z)|\leq 16$ holds for $x\in[-1/2,0]$ and $|y|\leq 10^{2}$. Finally, $|\chi(z)|\leq 1.1$ for $x\in[1/2,1]$ and $|y|\geq 2\pi$, while $|(z-1)\chi(z)|\leq 6.31$ for $x\in[1/2,1]$ and $|y|\leq 2\pi$ by the maximum principle and numerical verification. ∎
@@
$\left|\int_{\sigma_{3}+\mathrm{i}T}^{\sigma_{3}+2\mathrm{i}T}F_{2}(z)\mathrm{d}z\right|\leq\frac{1}{\left(\mathcal{U}^{-}(T_{0})\right)^{2}}\left(\frac{2\pi}{T}\right)^{1-2\sigma_{3}}\left(\int_{T}^{2T}\left|\sum_{n\leq T}\frac{d(n)}{n^{\sigma_{3}+\mathrm{i}u}}\right|^{2}\mathrm{d}u\right)^{\frac{1}{2}}\times\\ \times\left(\left(\int_{T}^{2T}\left|\sum_{n>T}\frac{d(n)}{n^{\sigma_{3}+\mathrm{i}u}}e^{-n/T}\right|^{2}\mathrm{d}u\right)^{\frac{1}{2}}+\left(\int_{T}^{2T}\left|\sum_{n\leq T}\frac{d(n)}{n^{\sigma_{3}+\mathrm{i}u}}\left(e^{-n/T}-1\right)\right|^{2}\mathrm{d}u\right)^{\frac{1}{2}}\right).$
by the Cauchy–Bunyakovsky–Schwarz (CBS) inequality and Lemma 18. Then
@@
×$\times\left(\left(\eta(T_{0})+2\pi m_{0}\right)^{\frac{1}{2}}\mathcal{D}_{7}(T_{0})^{\frac{1}{2}}+\left(\eta(T_{0})\mathcal{D}_{4}(T_{0},2\sigma_{2}-2)+2\pi m_{0}\mathcal{D}_{4}(T_{0},2\sigma_{2}-3)\right)^{\frac{1}{2}}\right)$
@@
×$\times\left(\left(\eta(T_{0})+2\pi m_{0}\right)^{\frac{1}{2}}\mathcal{D}_{7}(T_{0})^{\frac{1}{2}}+\left(\eta(T_{0})\mathcal{D}_{4}(T_{0},2\sigma_{3}-2)+2\pi m_{0}\mathcal{D}_{4}(T_{0},2\sigma_{3}-3)\right)^{\frac{1}{2}}\right).$ (81)
@@
We need to estimate the moduli of each of the above terms. Starting with the first term on the right-hand side of (84), note that $20<1-\sigma_{4}<1/$ 2. By the CBS inequality and Lemma 18 we then have
@@
We estimate the second integral on the right-hand side of (85). Note that $-1/2<c_{1}+\sigma_{4}<0$, implying $|c_{1}+\sigma_{4}+\mathrm{i}(t+u)|\leq 1/2+2T+\mathfrak{a}_{1}(\log{T})^{\mathfrak{a}_{2}}$ since $t\in[T,2T]$ and $|u|\leq\mathfrak{a}_{1}\left(\log{T}\right)^{\mathfrak{a}_{2}}$. By the CBS inequality and Lemma 18 we thus have
@@
×$\times\left(\mathcal{U}^{+}\left(T_{0}-\mathfrak{a}_{1}(\log{T_{0}})^{\mathfrak{a}_{2}}\right)\right)^{4}\int_{-\infty}^{\infty}\left|\Gamma(c_{1}+\mathrm{i}u)\right|\mathrm{d}u\times$
×$\times\int_{-\mathfrak{a}_{1}\left(\log{T}\right)^{\mathfrak{a}_{2}}}^{\mathfrak{a}_{1}\left(\log{T}\right)^{\mathfrak{a}_{2}}}\left|\Gamma(c_{1}+\mathrm{i}u)\right|\left(\int_{T}^{2T}\left|\sum_{n>T}\frac{d(n)}{n^{1-\sigma_{4}-c_{1}-\mathrm{i}(t+u)}}\right|^{2}\mathrm{d}t\right)\mathrm{d}u.$
@@
×$\times\left(\mathcal{U}^{+}\left(T_{0}-\mathfrak{a}_{1}(\log{T_{0}})^{\mathfrak{a}_{2}}\right)\right)^{4}\left(\int_{-\infty}^{\infty}\left|\Gamma(c_{1}+\mathrm{i}u)\right|\mathrm{d}u\right)^{2}\times$
×$\times\left(\frac{\eta(T_{0})\mathcal{D}_{5}\left(T_{0},2-2\sigma_{4}+2|c_{1}|\right)}{(2\pi)^{2}(1-2\sigma_{4}+2|c_{1}|)^{4}}+\frac{m_{0}\mathcal{D}_{5}\left(T_{0},1-2\sigma_{4}+2|c_{1}|\right)}{2\pi\left(2|c_{1}|-2\sigma_{4}\right)^{4}}\right).$ (87)
Here, $\mathcal{D}_{5}(x_{0},\sigma)$ and $\mathcal{U}^{\pm}(y_{0})$ are defined by (14) and (73), respectively. Turning to the third integral on the right-hand side of (85), note that 1/$21/2<c_{2}+\sigma_{4}<3/$ 2 and $|\sigma_{4}+c_{2}+\mathrm{i}(t+u)|\geq T-\mathfrak{b}_{1}(\log{T})^{\mathfrak{b}_{2}}$ since $t\in[T,2T]$ and $|u|\leq\mathfrak{b}_{1}\left(\log{T}\right)^{\mathfrak{b}_{2}}$. Also, $-1<2(1-\sigma_{4}-c_{2})<1$. Similarly as before, the CBS inequality and Lemmas 1014 and 18 imply
@@
×$\times\left(\mathcal{U}^{+}\left(T_{0}-\mathfrak{b}_{1}\left(\log{T_{0}}\right)^{\mathfrak{b}_{2}}\right)\right)^{4}\left(\int_{-\infty}^{\infty}\left|\Gamma(c_{2}+\mathrm{i}u)\right|\mathrm{d}u\right)^{2}\times$
×$\times\left(\frac{\eta(T_{0})\mathcal{D}_{4}\left(T_{0},2\left(1-\sigma_{4}-c_{2}\right)\right)}{(2\pi)^{2}}+\frac{m_{0}\mathcal{D}_{4}\left(T_{0},1-2\left(\sigma_{4}+c_{2}\right)\right)}{2\pi}\right).$ (89)
@@
×$\times\left(\mathfrak{g}_{1}\left(T_{0},\sigma_{4},c_{1},\bm{\mathfrak{a}}\right)^{\frac{1}{2}}+\mathfrak{g}_{2}\left(T_{0},\sigma_{4},c_{2},\bm{\mathfrak{b}}\right)^{\frac{1}{2}}\right).$ (91)
@@
by the CBS inequality and Lemma 18 since $20<\sigma_{5}<1/$ 2. Estimation of the second and the third integral is similar as before. We obtain
@@
×$\times\left(\mathcal{U}^{+}\left(T_{0}-\mathfrak{a}_{1}(\log{T_{0}})^{\mathfrak{a}_{2}}\right)\right)^{4}\left(\int_{-\infty}^{\infty}\left|\Gamma(c_{1}+\mathrm{i}u)\right|\mathrm{d}u\right)^{2}\times$
×$\times\left(\frac{\eta(T_{0})\mathcal{D}_{5}(T_{0},2-2\sigma_{5}+2|c_{1}|)}{\left(1-2\sigma_{5}+2|c_{1}|\right)^{4}}+\frac{2\pi m_{0}\mathcal{D}_{5}(T_{0},1-2\sigma_{5}+2|c_{1}|)}{\left(2|c_{1}|-2\sigma_{5}\right)^{4}}\right)$ (92)
@@
×$\times\left(\mathcal{U}^{+}\left(T_{0}-\mathfrak{b}_{1}\left(\log{T_{0}}\right)^{\mathfrak{b}_{2}}\right)\right)^{4}\left(\int_{-\infty}^{\infty}\left|\Gamma(c_{2}+\mathrm{i}u)\right|\mathrm{d}u\right)^{2}\times$
×$\times\left(\eta(T_{0})\mathcal{D}_{4}\left(T_{0},2\left(1-\sigma_{5}-c_{2}\right)\right)+2\pi m_{0}\mathcal{D}_{4}\left(T_{0},1-2\left(\sigma_{5}+c_{2}\right)\right)\right).$ (93)

### v2

@@ 4. Proof of Theorem 1
$F_{X}\left(\sigma,H,T,w\right)\leq\left(\frac{\exp{\left(\kappa\left(1+\left(\frac{1}{4T}\left(1+\frac{2\delta}{\log{(hT)}}\right)\right)^{2}\right)\right)}}{1-H^{-1}}\right)^{\frac{1}{w}}\times\\ \times\left(\left(K_{1}^{\frac{\frac{3}{4}w_{1}}{w}}K_{3}^{\frac{\frac{1}{2}w_{2}}{w}}\right)T^{\frac{\frac{3}{4}w_{1}}{w}}\left(\log{T}\right)^{\frac{w_{1}}{w}}\left(\log{(hT)}\right)^{\frac{\frac{1}{2}\left(w_{1}+w_{2}\right)}{w}}+\left(K_{2}^{\frac{\frac{3}{4}w_{1}}{w}}K_{3}^{\frac{\frac{1}{2}w_{2}}{w}}\right)T^{\frac{\frac{3}{4}w_{1}}{w}}\left(\log{(hT)}\right)^{\frac{\frac{1}{2}w_{2}}{w}}\right)$ (42)
@@
$K_{3}^{\frac{\frac{1}{2}w_{2}}{w}}\leq\max\left\{1,K_{3}^{-\frac{3\delta}{\log{(hT_{0})}}}\right\}K_{3}^{\frac{2\sigma-1}{2-\sigma}},\quad\left(\log{(hT)}\right)^{\frac{\frac{1}{2}w_{2}}{w}}\leq\left(\log{(hT)}\right)^{\frac{2\sigma-1}{2-\sigma}}$ (45)
since $\log{(hT)}\geq 1$. Turning to the other log-terms, because $\log{T}\geq 1$ and
@@
$\left(\log{T}\right)^{\frac{w_{1}}{w}}\leq\exp{\left(\frac{4\delta\log{\log{T}}}{\log{(hT)}}\right)}\left(\log{T}\right)^{\frac{4(1-\sigma)}{2-\sigma}}\leq\exp{\left(4\delta\max\left\{\frac{\log{\log{T_{0}}}}{\log{(hT_{0})}},\frac{\log{\log{T_{0}}}}{\log{T_{0}}}\right\}\right)}\left(\log{T}\right)^{\frac{4(1-\sigma)}{2-\sigma}},$ (46)
where we also used $T_{0}\geq e^{e}$. Moreover, $w_{1}+w_{2}=1$ and
$\frac{1}{w}=\frac{2+\frac{4\delta}{\log{(hT)}}}{2-\sigma+\frac{3\delta}{\log{(hT)}}}\leq\frac{2}{2-\sigma},$ (47)
@@
$\left(\log{(hT)}\right)^{\frac{\frac{1}{2}\left(w_{1}+w_{2}\right)}{w}}\leq\left(\log{(hT)}\right)^{\frac{1}{2-\sigma}}.$ (48)
In addition, estimate (47) is used to bound the first term on the right-hand side of (42), and this contributes $L_{0}$ from (35) to the final estimate. Inequalities (43)–(46) and (48) contribute to $L_{1}$ from (39), while inequalities (43) (for $K_{2}$ in place of $K_{1}$), (44) and (45), contribute to $L_{2}$ from (40). The proof of Lemma 9 is thus complete. ∎
@@
$\times\left(L_{1}\max\left\{1,K_{1}^{\frac{3d}{\log{T_{0}}}}\right\}\exp{\left(d\left(3+\frac{4\log{\log{T_{0}}}}{\log{T_{0}}}\right)\right)}\left(K_{1}T\right)^{\frac{3(1-\sigma)}{2-\sigma}}\left(\log{T}\right)^{\frac{4(1-\sigma)}{2-\sigma}}\left(\log{(hT)}\right)^{\frac{1}{2-\sigma}}\right$.
$\left.+L_{2}e^{3d}\max\left\{1,K_{2}^{\frac{3d}{\log{T_{0}}}}\right\}\left(K_{2}T\right)^{\frac{3(1-\sigma)}{2-\sigma}}\left(\log{(hT)}\right)^{\frac{2\sigma-1}{2-\sigma}}\right),$
where $L_{0}$, $L_{1}$, $L_{2}$ and $K_{1}$, $K_{2}$, $K_{3}$ are as in Lemma 9. It is not hard to see that the statement of Theorem 1 now follows after combining all estimates. Also, the non-vanishing condition can be removed by invoking the standard perturbation argument since the right-hand side of the main estimate of Theorem 1 is a continuous function in $\sigma$, $T$ and $H$. ∎
@@ 5.1. The method
This is a meromorphic function on $\mathbb{C}$ with only simple poles at s=2k1s=2k-1 for $k\in\mathbb{N}$. Also note that the functional equation implies $\chi(1/2+\mathrm{i}t)\chi(1/2-\mathrm{i}t)=\left|\chi(1/2\pm\mathrm{i}t)\right|^{2}=1$. The starting point in Ramachandra’s proof (which itself resembles the reflection principle [Ivi03, Section 4.4] or a half-way derivation of a weighted approximate functional equation [Ivi13, Section 4.5]) is to use the Mellin integral for $e^{-x}$ to get
@@
The curly brackets in (54) and (55) indicate the plan of the proof of Theorem 2. Precisely, the seemingly unnatural appearance of the second and the third part allows us to obtain an asymptotically correct estimate (12) since we can show that all terms, except the first one, are $\ll T\log^{3}{T}$. However, if one were interested in obtaining just an upper bound like (13), then  (54) and (55) can be simplified to
@@ 5.2. Some results on various sums that contain $d(n)$ and $d^{2}(n)$
Remark 2.
Although there exist better explicit estimates for $|\Delta(x)|$, e.g., of the form $\Delta(x)\ll x^{\frac{1}{3}}\log x$ as in [BBR12], the above weaker version serves our purpose.
@@
by partial summation and Lemma 13. Repeated use of integration by parts gives
@@
$e^{-2}x^{2}\sum_{k=1}^{4}D_{k}(\log{x})^{4-k}+\sum_{k=1}^{4}D_{k}\int_{x}^{\infty}\left(\frac{2}{x}u^{2}-u\right)e^{-\frac{2u}{x}}(\log{u})^{4-k}\mathrm{d}u\leq\frac{3D_{1}}{4e^{2}}x^{2}(\log{x})^{3}\\ +\left(\frac{57D_{1}}{16e^{2}}+\frac{3D_{2}}{4e^{2}}\right)x^{2}(\log{x})^{2}+\left(\frac{3D_{1}}{e^{2}}+\frac{19D_{2}}{8e^{2}}+\frac{3D_{3}}{4e^{2}}\right)x^{2}\log{x}+\left(\frac{3D_{1}}{4e^{2}}+\frac{D_{2}}{e^{2}}+\frac{19D_{3}}{16e^{2}}+\frac{3D_{4}}{4e^{2}}\right)x^{2}.$ (73)
@@ 5.3. Proof of Theorem 2
with the last inequalities true by [DZA22, Proposition 2.2]. For the second part, one can use (75) to obtain $|\chi(z)|\leq 0.4|y|^{1/2-x}$ for $x\in[-1/2,0]$ and $|y|\geq 10^{2}$, while the maximum principle and numerical verification show that $|\chi(z)|\leq 16$ holds for $x\in[-1/2,0]$ and $|y|\leq 10^{2}$. Finally, $|\chi(z)|\leq 1.1$ for $x\in[1/2,1]$ and $|y|\geq 2\pi$, while $|(z-1)\chi(z)|\leq 6.31$ for $x\in[1/2,1]$ and $|y|\leq 2\pi$ by the maximum modulus principle and numerical verification. ∎
@@
$\left|\int_{\sigma_{3}+\mathrm{i}T}^{\sigma_{3}+2\mathrm{i}T}F_{2}(z)\mathrm{d}z\right|\leq\frac{1}{\left(\mathcal{U}^{-}(T_{0})\right)^{2}}\left(\frac{2\pi}{T}\right)^{1-2\sigma_{3}}\left(\int_{T}^{2T}\left|\sum_{n\leq T}\frac{d(n)}{n^{\sigma_{3}+\mathrm{i}u}}\right|^{2}\mathrm{d}u\right)^{\frac{1}{2}}\times\\ \times\left(\left(\int_{T}^{2T}\left|\sum_{n>T}\frac{d(n)}{n^{\sigma_{3}+\mathrm{i}u}}e^{-n/T}\right|^{2}\mathrm{d}u\right)^{\frac{1}{2}}+\left(\int_{T}^{2T}\left|\sum_{n\leq T}\frac{d(n)}{n^{\sigma_{3}+\mathrm{i}u}}\left(e^{-n/T}-1\right)\right|^{2}\mathrm{d}u\right)^{\frac{1}{2}}\right)$
by the CBS inequality and Lemma 18. Then
@@
$\times\left(\left(\eta(T_{0})+2\pi m_{0}\right)^{\frac{1}{2}}\mathcal{D}_{7}(T_{0})^{\frac{1}{2}}+\left(\eta(T_{0})\mathcal{D}_{4}(T_{0},2\sigma_{2}-2)+2\pi m_{0}\mathcal{D}_{4}(T_{0},2\sigma_{2}-3)\right)^{\frac{1}{2}}\right)$
@@
$\times\left(\left(\eta(T_{0})+2\pi m_{0}\right)^{\frac{1}{2}}\mathcal{D}_{7}(T_{0})^{\frac{1}{2}}+\left(\eta(T_{0})\mathcal{D}_{4}(T_{0},2\sigma_{3}-2)+2\pi m_{0}\mathcal{D}_{4}(T_{0},2\sigma_{3}-3)\right)^{\frac{1}{2}}\right).$ (83)
@@
We need to estimate the moduli of each of the above terms. Starting with the first term on the right-hand side of (86), note that $0<1-\sigma_{4}<1/2$. By the CBS inequality and Lemma 18 we then have
@@
We estimate the second integral on the right-hand side of (87). Note that $1/2<c_{1}+\sigma_{4}<0$, implying $|c_{1}+\sigma_{4}+\mathrm{i}(t+u)|\leq 1/2+2T+\mathfrak{a}_{1}(\log{T})^{\mathfrak{a}_{2}}$ since $t\in[T,2T]$ and $|u|\leq\mathfrak{a}_{1}\left(\log{T}\right)^{\mathfrak{a}_{2}}$. By the CBS inequality and Lemma 18 we thus have
@@
$\times\left(\mathcal{U}^{+}\left(T_{0}-\mathfrak{a}_{1}(\log{T_{0}})^{\mathfrak{a}_{2}}\right)\right)^{4}\int_{-\infty}^{\infty}\left|\Gamma(c_{1}+\mathrm{i}u)\right|\mathrm{d}u\times$
$\times\int_{-\mathfrak{a}_{1}\left(\log{T}\right)^{\mathfrak{a}_{2}}}^{\mathfrak{a}_{1}\left(\log{T}\right)^{\mathfrak{a}_{2}}}\left|\Gamma(c_{1}+\mathrm{i}u)\right|\left(\int_{T}^{2T}\left|\sum_{n>T}\frac{d(n)}{n^{1-\sigma_{4}-c_{1}-\mathrm{i}(t+u)}}\right|^{2}\mathrm{d}t\right)\mathrm{d}u.$
@@
$\times\left(\mathcal{U}^{+}\left(T_{0}-\mathfrak{a}_{1}(\log{T_{0}})^{\mathfrak{a}_{2}}\right)\right)^{4}\left(\int_{-\infty}^{\infty}\left|\Gamma(c_{1}+\mathrm{i}u)\right|\mathrm{d}u\right)^{2}\times$
$\times\left(\frac{\eta(T_{0})\mathcal{D}_{5}\left(T_{0},2-2\sigma_{4}+2|c_{1}|\right)}{(2\pi)^{2}(1-2\sigma_{4}+2|c_{1}|)^{4}}+\frac{m_{0}\mathcal{D}_{5}\left(T_{0},1-2\sigma_{4}+2|c_{1}|\right)}{2\pi\left(2|c_{1}|-2\sigma_{4}\right)^{4}}\right).$ (89)
Here, $\mathcal{D}_{5}(x_{0},\sigma)$ and $\mathcal{U}^{\pm}(y_{0})$ are defined by (14) and (75), respectively. Turning to the third integral on the right-hand side of (87), note that $1/2<c_{2}+\sigma_{4}<3/2$ and $|\sigma_{4}+c_{2}+\mathrm{i}(t+u)|\geq T-\mathfrak{b}_{1}(\log{T})^{\mathfrak{b}_{2}}$ since $t\in[T,2T]$ and $|u|\leq\mathfrak{b}_{1}\left(\log{T}\right)^{\mathfrak{b}_{2}}$. Also, $1<2(1-\sigma_{4}-c_{2})<1$. Similarly as before, the CBS inequality and Lemmas 1014 and 18 imply
@@
$\times\left(\mathcal{U}^{+}\left(T_{0}-\mathfrak{b}_{1}\left(\log{T_{0}}\right)^{\mathfrak{b}_{2}}\right)\right)^{4}\left(\int_{-\infty}^{\infty}\left|\Gamma(c_{2}+\mathrm{i}u)\right|\mathrm{d}u\right)^{2}\times$
$\times\left(\frac{\eta(T_{0})\mathcal{D}_{4}\left(T_{0},2\left(1-\sigma_{4}-c_{2}\right)\right)}{(2\pi)^{2}}+\frac{m_{0}\mathcal{D}_{4}\left(T_{0},1-2\left(\sigma_{4}+c_{2}\right)\right)}{2\pi}\right).$ (91)
@@
$\times\left(\mathfrak{g}_{1}\left(T_{0},\sigma_{4},c_{1},\bm{\mathfrak{a}}\right)^{\frac{1}{2}}+\mathfrak{g}_{2}\left(T_{0},\sigma_{4},c_{2},\bm{\mathfrak{b}}\right)^{\frac{1}{2}}\right).$ (93)
@@
by the CBS inequality and Lemma 18 since $0<\sigma_{5}<1/2$. Estimation of the second and the third integral is similar as before. We obtain
@@
$\times\left(\mathcal{U}^{+}\left(T_{0}-\mathfrak{a}_{1}(\log{T_{0}})^{\mathfrak{a}_{2}}\right)\right)^{4}\left(\int_{-\infty}^{\infty}\left|\Gamma(c_{1}+\mathrm{i}u)\right|\mathrm{d}u\right)^{2}\times$
$\times\left(\frac{\eta(T_{0})\mathcal{D}_{5}(T_{0},2-2\sigma_{5}+2|c_{1}|)}{\left(1-2\sigma_{5}+2|c_{1}|\right)^{4}}+\frac{2\pi m_{0}\mathcal{D}_{5}(T_{0},1-2\sigma_{5}+2|c_{1}|)}{\left(2|c_{1}|-2\sigma_{5}\right)^{4}}\right)$ (94)
@@
$\times\left(\mathcal{U}^{+}\left(T_{0}-\mathfrak{b}_{1}\left(\log{T_{0}}\right)^{\mathfrak{b}_{2}}\right)\right)^{4}\left(\int_{-\infty}^{\infty}\left|\Gamma(c_{2}+\mathrm{i}u)\right|\mathrm{d}u\right)^{2}\times$
$\times\left(\eta(T_{0})\mathcal{D}_{4}\left(T_{0},2\left(1-\sigma_{5}-c_{2}\right)\right)+2\pi m_{0}\mathcal{D}_{4}\left(T_{0},1-2\left(\sigma_{5}+c_{2}\right)\right)\right).$ (95)